OpenStudy (anonymous):

Write an equation in point-slope form, y – y1 = m(x – x1), of the line through points (9, –2) and (10, 2). Use (9, –2) as the point (x1, y1).

OpenStudy (anonymous):

\[slop=\frac{ (y-y_1)}{(x-x_1)}\] \[slop=\frac{2-(-2)}{10-9}=\frac{ 4 }{ 1 }=4\] So the equation of the line with regards to those two points is: \[y-y_1=slop*(x-x_1)\] \[y-(-2)=4*(x-9)\] \[y+2=4x-36\] Thus: \[y=4x-38\]